OpenStudy (anonymous):

what is teh slope of 3x+4y=1

OpenStudy (zzr0ck3r):

hi

OpenStudy (zzr0ck3r):

you can close this

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how is ur cat

OpenStudy (zzr0ck3r):

hes fine. Just had to get boosters

OpenStudy (zzr0ck3r):

wife got stuck at work

OpenStudy (anonymous):

i see. ah. so u had to take him. ur a good cat dad.

OpenStudy (zzr0ck3r):

you have no idea

OpenStudy (zzr0ck3r):

lol

OpenStudy (anonymous):

thanks for helping out. the hw over the weekend almost made me want to quit math. lol

OpenStudy (zzr0ck3r):

lol

OpenStudy (anonymous):

but i feel better about it now that everyone else struggled too

OpenStudy (zzr0ck3r):

that's how it should be:)

OpenStudy (zzr0ck3r):

what you working on?

OpenStudy (anonymous):

why?? shouldn't math be roses and cotton candy??

OpenStudy (zzr0ck3r):

:)

OpenStudy (anonymous):

cptr 10...orders...the proofs. i cant' seem to find the logic to create a proof

OpenStudy (zzr0ck3r):

yeah that's a tough section

OpenStudy (anonymous):

not to create...but to make the proof work out

OpenStudy (anonymous):

do u have the book?

OpenStudy (zzr0ck3r):

yeah

OpenStudy (anonymous):

E2. i tried it over and over but couldn't make something happen

OpenStudy (zzr0ck3r):

m and n relative prime one?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

i don't know how to integrate divisors with orders

OpenStudy (anonymous):

there is a disconnect in my mind of how the two relate

OpenStudy (zzr0ck3r):

give me a few to think about it

OpenStudy (anonymous):

k.

OpenStudy (anonymous):

Let 0 < i < m และ 0 < j < n suppose ai = bj some i,j e = am = ei = (am)i = ami = (ai)m = (bj)m =bjm so n | mj but gcd(m,n) = 1 n | j It's conflict because j<n ai bj for all integer i,j which not 0

OpenStudy (anonymous):

its chinese to me

OpenStudy (zzr0ck3r):

ok

OpenStudy (zzr0ck3r):

assume by contradiction there is a j<m and k<n such that a^j = a^k the reason j<m is because every thing after a^m is just a repeat of some multiple of a, call it a^j

OpenStudy (zzr0ck3r):

does this make sense?

OpenStudy (zzr0ck3r):

if there is a v > m st a^v is defined then that same element is a^j for some j < m

OpenStudy (anonymous):

so m is the smallest positive integer?

OpenStudy (zzr0ck3r):

smallest positive intiger st a^m = e

OpenStudy (anonymous):

ok

OpenStudy (zzr0ck3r):

the only important thing is to understand why we can say j<m and K < n

OpenStudy (zzr0ck3r):

its just because if they were greater, we could find that same element before with hit e

OpenStudy (zzr0ck3r):

say a^3 = e we have a,a^1,a^2,a^3,a^4 note that a^4 = a^1

OpenStudy (anonymous):

yes, i understand that

OpenStudy (zzr0ck3r):

ok pellet, my dad is here I just have to drop him off about 4 blocks away. can you stay?

OpenStudy (zzr0ck3r):

10 mins at most

OpenStudy (anonymous):

yes, no worries. i will be here.

OpenStudy (anonymous):

take ur time

OpenStudy (zzr0ck3r):

still here

OpenStudy (anonymous):

yes. i'm here. i reviewed what u wrote...and it makes sense so far

OpenStudy (zzr0ck3r):

assume by contradiction there is a j<m and k<n such that a^j = a^k \[e=a^m\] since m is the order of a, we know that a raised to any multiple of m is also = e \[e=a^m = a^{mi}\]

OpenStudy (anonymous):

k

OpenStudy (zzr0ck3r):

\[e = a^m = a^{mi}= (a^i)^m\] since a^i = b^j we have \[e = a^m = a^{mi}= (a^i)^m = (b^j)^m\]

OpenStudy (zzr0ck3r):

so assume by contradiction that a^i = b^j \[e= a^m = a^{mi}= (a^i)^m= (b^j)^m\]

OpenStudy (zzr0ck3r):

sorry about that

OpenStudy (zzr0ck3r):

this make sense?

OpenStudy (anonymous):

ok. maybe a stupid question...when was it established that a^i = b^j

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

its a contradiction

OpenStudy (zzr0ck3r):

the thing we are trying to prove is that if ord(a) = m, and ord(b) = n then a^i=b^j cant be possible when (a,m) = 1 so yes we assume by contradiction that a^i = b^j for some I,j

OpenStudy (anonymous):

(n,m)=1?

OpenStudy (anonymous):

i get it...but it just seems to far fetched for us group theory students to figure out on our own without proper training.

OpenStudy (anonymous):

*too

OpenStudy (zzr0ck3r):

(n,m) = 1 just means n,m are relative prime

OpenStudy (zzr0ck3r):

ill show you when we are done the login into coming up with it on your own

OpenStudy (anonymous):

yes, just confirming (n,m)=1 not (a,m)=1

OpenStudy (zzr0ck3r):

o yeah:)

OpenStudy (zzr0ck3r):

so far we have PF: let G be a group let a,b be in G s.t. ord(a) = m and ord(b) = n where n,m are relatively prime assume by contradiction there exists I,j in Z s.t. a^i = b^j then there is i<m and j<n st a^i - b^j then \[e=a^m = a^{mi} = (a^i)^m = (b^j)^m\]

OpenStudy (zzr0ck3r):

\[(b^j)^m = b^{jm}\] so we have \[e = b^{jm}\]

OpenStudy (zzr0ck3r):

right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

got it

OpenStudy (zzr0ck3r):

since n is the order of b, n is the smallest number such that b^n = e so we know that jm is a multiple of n because b^(jm) = e so n divides jm n|jm we know n does not divide m because they are relative prime, and we know n does not divide j because n<j, so we have a contradiction that n|jm

OpenStudy (zzr0ck3r):

so far we have PF: let G be a group let a,b be in G s.t. ord(a) = m and ord(b) = n where n,m are relatively prime assume by contradiction there exists I,j in Z s.t. a^i = b^j then there is i<m and j<n st a^i - b^j then \[e=a^m = a^{mi} = (a^i)^m = (b^j)^m = b^{jm}\] that is jm is a multiple of ord(b) = n so n|mj but (n,m) = 1 and j<n by assumption this is a contradiction and no such i,j exist. qed.

OpenStudy (anonymous):

ok. i follow the logic...but most of the kids in my class couldn't figure this out. i feel like it requires more proof experience...as we have not learned by contradiction yet

OpenStudy (zzr0ck3r):

as far as trying to come up with this. first thing you ask yourself is what do I need to show? well we need to show that a^i = b^j for some I,j given the conditions we have how do we show this is true^^^^^ THE HELL IF I KNOW this screams contradiction. The fact that I cant even figure out how to say what we need to show screams contradiction because how do we show something for all I,j without induction ( im sure there is an inductive proof).

OpenStudy (anonymous):

gerardo assigned it adn it was one of the problems due to day. i did the best i could but eventually just wrote him a short note that i was stuck with a frowny face

OpenStudy (anonymous):

ah. ok. so its a clue that i may have to use contradiction when i am pulling my hair out

OpenStudy (zzr0ck3r):

then we say ok how do we start the contradiction proof assume the thing a^i = b^j and write one thing you know for sure a^m = e (note we could have started with b^n = e and done the same thing)

OpenStudy (zzr0ck3r):

so far we have had to come up with nothing.. so here is the tricky part a^(im) = e once we have that, the proof is all natural

OpenStudy (anonymous):

yeah..that is how i started...but never finished. its ok. its a learning experience

OpenStudy (zzr0ck3r):

yeah for sure

OpenStudy (anonymous):

thank u for ur help...gotta run...husband getting off of work soon.

OpenStudy (zzr0ck3r):

this order stuff has the most ambiguous proofs, they are odd but really they are all alike. I think if you understand this one well. you will feel much better with the others

OpenStudy (zzr0ck3r):

np

OpenStudy (zzr0ck3r):

ill be around later whenever you want. just text

OpenStudy (anonymous):

i will redo it over teh next couple of days

OpenStudy (anonymous):

i have an exam in grp thry on thursday

OpenStudy (zzr0ck3r):

ill show you the other "method" for order proofs with m = qn+r

OpenStudy (anonymous):

thx zach...much appreciated!!

OpenStudy (zzr0ck3r):

np sorry about today

OpenStudy (anonymous):

don't worry...not a big deal. talk to u soon!!

OpenStudy (zzr0ck3r):

bye